# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ Abstract.
Many intended uses of differential privacy involve a continual mechanism that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated. In this paper, we give the first general treatment of privacy against adaptive adversaries for mechanisms that support dataset updates and a variety of queries, all arbitrarily interleaved. It also models a very general notion of neighboring, that includes both event-level and user-level privacy.
We prove several concurrent composition theorems for continual mechanisms, which ensure privacy even when an adversary can interleave its queries and dataset updates to the different composed mechanisms. Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates. Moreover, we also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy. Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates.
@@ 1. Introduction
[…]
@@ 1.1. Differential Privacy Over Time
Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature. For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters were given by [19]. Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 11, 17] and the Private Multiplicative Weights algorithm [11].
@@
A few years ago, Vadhan and Wang [20] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries. Specifically, they formalized the notion of an interactive mechanism, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time. (Both SVT and Private Multiplicative Weights utilize a secret state.) They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
@@
Vadhan and Wang [20] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms. In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed. This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks. Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case. A tight bound for concurrent composition of approximate differential privacy was given by Lyu [18]. A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [18], $f$-DP [21], and adaptive choices of privacy parameters [10].
@@
All of the above works are for the case where the queries to the mechanisms are adaptive, but the dataset $x$ is static, given as input to the mechanism at the start of the interaction. Jain, Raskhodnikova, Sivakumar, and Smith [13] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update. The lack of a more general formalism and toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g. [12]). Rectifying this situation is what motivated our work.
@@ 1.2. Our contributions
In this paper, we:
  • Give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates. This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries). See Figure 1 for an illustration. We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries [14], but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
@@
  • We prove concurrent composition theorems for continual mechanisms. (1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms. (2) We extend the “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of a finite or infinite number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets. (3) Then we extend this further to the concurrent composition of a finite or infinite number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy. (4) We complement the latter with a counterexample for the case of approximate differential privacy.
  • We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss.
  • We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [12]. In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
@@
For every fixed sequence $(\epsilon_{1},\delta_{1}),\ldots,(\epsilon_{k},\delta_{k})$ of privacy-loss parameters and verification functions $f_{1},\ldots,f_{k}$, the concurrent composition of continual mechanisms $\mathcal{M}_{i}$ that are $(\epsilon_{i},\delta_{i})$-DP wrt $f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
@@
Theorem 1.2 is proven by using the aforementioned reduction from continual to interactive mechanism. Then we use Lyu’s result  [18] for $(\epsilon,\delta)$-DP interactive mechanisms and extend it to hold for infinite (but discrete) answer spaces and an unbounded number of rounds of interaction as follows. In Lyu’s proof the restriction on the number of interactions comes from a backward definition of a so-called control function. We define a function which is the limit over this control function and, using monotone convergence, show that this “limit function” satisfies all the necessary conditions to be used in the rest of Lyu’s proof. This extends the result to an unbounded number of rounds of interaction. Furthermore, Lyu’s proof used an algorithmic construction iterating over the answer set, implying that the answer set must be finite. Instead with a mathematical approach, we extend the proof to work with a countable, infinite answer set.
@@
We study the extent to which parallel composition extends to the concurrent composition of continual mechanisms, meaning the neighboring relation is as above, but the adversary can interact concurrently with the mechanisms, and in particular decide adaptively on which $k$ of them to have different update sequences. Moreover, we also consider the case where $\ell$ is unbounded, meaning the adversary (or honest analyst) can adaptively decide how many mechanisms to interact with. We refer to this as concurrent $k$-sparse parallel composition.
@@
For every $\epsilon_{0},\delta_{0}\geq 0$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
As discussed later in the introduction, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
@@
Theorem 1.4 (concurrent parallel composition of approximate DP continual mechanisms, informally stated).
@@
This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms.
However, we prove that if $\delta_{0}=0$, concurrent parallel composition behaves just like non-interactive parallel composition:
@@
Theorem 1.5 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
@@
For every $\epsilon_{0}\geq 0$ and integers $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
We note that the constraint of pure DP is only on the mechanisms being composed; we allow $\delta>0$ in the final privacy guarantee, so we can still, for example, recover the advanced composition bound applied to pure DP mechanisms. Like Theorem 1.3, we also prove a generalization of Theorem 1.5 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
To prove this theorem, we add a twist to the concurrent composition theorems for interactive mechanisms. Specifically, the papers [20, 18] proved their composition theorems for pure and approximate DP by showing that for every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}$ and pair $x_{0}\sim x_{1}$ of adjacent datasets, $\mathcal{M}$ can be simulated as an interactive post-processing of the noninteractive $(\epsilon_{0},\delta_{0})$-DP randomized response mechanism $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$. That is, there is an IPM $\mathcal{T}$ (with no input) such that for each $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is equivalent to $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon_{0},\delta_{0}}(b)$.
We show that when $\delta_{0}=0$, the use of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ by $\mathcal{T}$ can be delayed until the time that it receives a query that would have different answer distributions under $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$. While intuitive, Theorem 1.4 tells us that this fact fails when $\delta_{0}>0$.
It is an interesting open problem to see if Theorem 1.5 can be extended to composing approximate DP mechanisms ($\delta_{0}>0$), if we allow an additional slackness in the final $\delta$, as shown necessary by Theorem 1.4.
An Algorithmic Application.

### v3

@@ Abstract.
Many intended uses of differential privacy involve a continual mechanism that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated. In this paper, we give the first general treatment of privacy against adaptive adversaries for mechanisms that support dataset updates and a variety of queries, all arbitrarily interleaved. It also models a very general notion of neighboring, that includes both event-level and user-level privacy. We prove several concurrent composition theorems for continual mechanisms, which ensure privacy even when an adversary can interleave its queries and dataset updates to the different composed mechanisms. Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates. We also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy. Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates. Thus, we prove a tight new composition theorem for this case. In addition, we prove concurrent filter compositions theorems for the scenarios in which the privacy parameters are adaptively chosen.
@@ 1. Introduction
Typically we think of $\epsilon$ as a small constant, e.g. $\epsilon=.1$, while $\delta$ is cryptographically negligible. The case that $\delta=0$ is referred to as pure differential privacy, and otherwise Inequality 1 is known as approximate differential privacy. This simple definition is very useful and easy to work with, but does not capture many scenarios in which we want to apply differential privacy.
@@ 1.1. Differential Privacy Over Time
Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature. For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by [21, 25]. Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 12, 18] and the Private Multiplicative Weights (PMW) algorithm [12].
@@
A few years ago, Vadhan and Wang [23] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries. Specifically, they formalized the notion of an interactive mechanism $\mathcal{M}$, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time. The use of a secret state, as utilized in both SVT and PMW, allows for correlating the answers to queries over time, and often allows for the privacy-loss budget to degrade much more slowly than composing noninteractive mechanisms would allow. They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
@@
Vadhan and Wang [23] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms. In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed. This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks.
The challenge in analyzing concurrent composition comes from the fact that the adversary can base its queries to $\mathcal{M}_{1}$ on answers given by $\mathcal{M}_{2}$ and vice-versa. The privacy guarantees on an interactive mechanism $\mathcal{M}$ only promise privacy against adversaries that do not get additional information about the sensitive dataset in the middle of its interaction with $\mathcal{M}$. Reasoning about the effect of such interleaved interactions is particularly tricky when the mechanisms $\mathcal{M}_{1}$ and/or $\mathcal{M}_{2}$ are not pure DP and can maintain secret state (like in the Sparse Vector Technique). Thus, it took several years before we had concurrent composition theorems that matched our composition theorems for noninteractive mechanisms.
Note that the extension of composition theorems for noninteractive mechanism to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive if it is differentially private against an adversary that was only interactive with $M$. Consider an adversary that first interacts adaptively with $M_{1}$, then $M_{2}$ and then with $M_{1}$ again. Note that for the second sequence of interactions with $M_{1}$ the adversary takes into account its prior interaction with $M_{1}$ as well as with $M_{2}$. Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive does not mean that it is adaptive against this interleaved composition of $M_{1}$ and $M_{2}$.
Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case. A tight bound for concurrent composition of approximate differential privacy was given by Lyu [19]. A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [19], $f$-DP [24], and adaptive choices of privacy parameters [11].
Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu [11] study the concurrent extension of composition theorems for non-interactive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters. This setting is referred to as filter composition, where a filter is a function $\mathrm{Filt}$ that maps the privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms to an element of $\{\top,\bot\}$. Suppose an adversary adaptively creates noninteractive mechanisms such that $g$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time. The $\mathrm{Filt}$-filter composition of noninteractive mechanisms is said to satisfy $(\epsilon,\delta)$-DP if the view of every such adversary is $(\epsilon,\delta)$-indistinguishable when all created mechanisms are run on dataset $x$ or its neighboring dataset $x^{\prime}$. Haney et al. extend this definition to the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms, where the adversary may not only create new mechanisms but also interact with existing interactive ones. They prove that if the $\mathrm{Filt}$-filter composition of non-interactive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
All of the above works are for the case where the queries to the mechanisms are adaptive, but the datasets are static, given as input to the mechanisms at the start of the interaction. [^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem [9], an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed. Thus, whenever $(x_{i+1},x_{i+1}^{\prime})$ are given, a new interactive mechanism $\mathcal{M}_{i+1}$ has to be started that does not know the secret state of $\mathcal{M}_{i}$. However, in some real-world applications, both queries and dataset updates may be issued adaptively, i.e., the dynamic updates to the dataset might be chosen adaptively. For instance, Ghazi et al. [10] study an advertising problem in which the server receives both private data and queries over time. They design algorithms that accept dataset updates while applying non-interactive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing non-interactive mechanisms with adaptive privacy-loss parameters. This independent work [^3] The first version of this work was available on arXiv prior to [10]. shows that if a mechanism receiving both queries and dataset updates produces a differentially private output at each time step $i$ (with adaptive parameters $\epsilon_{i},\delta_{i}$ at step $i$) such that $\sum_{i}\epsilon_{i}\leq\epsilon$ and $\sum_{i}\delta_{i}\leq\delta$, then the overall mechanism is $(\epsilon,\delta)$-DP. This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a non-interactive mechanism corresponding to a single step. While some continual mechanisms do indeed employ independent non-interactive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time. For these mechanisms, privacy must be analyzed over the entire output sequence rather than on a step-by-step basis to get the desired guarantees. (If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step non-interactive mechanisms. Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish [23, 19, 24, 11]. Thus, the results of Ghazi et al. [10] do not apply to the more general setting of (filter) concurrent composition of continual mechanisms. [^4] It is Ghazi et al.’s definition of Individual DP (IDP, their Definition 4.1) and their composition theorem that is framed in terms of IDP that restricts to step-by-step privacy analyses corresponding to composition of noninteractive mechanisms. Their definition of $(\epsilon,\delta)$-DP interactive mechanisms (their Definition 3.6) is somewhat more general than their definition of IDP, but it still does not allow for exploiting correlated randomness across queries to reduce privacy loss, as in SVT. (This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.
@@
Jain, Raskhodnikova, Sivakumar, and Smith [14] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update. The lack of a more general formalism and toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g. [13]). Rectifying this situation is what motivated our work.
Comparison with algorithms and data structures. We note that the different flavors of mechanisms discussed above are differential privacy analogues of familiar variants of randomized algorithms and data structures. A non-interactive mechanism is simply a static (or batch) algorithm. An interactive mechanism corresponds to static data structure. Our focus in this paper is on continual mechanisms, which allow both queries and updates, and these can be viewed as dynamic data structures. Adaptivity is a common concern for the correctness properties of randomized data structures; here we are concerned instead with its effect on privacy properties.
@@ 1.2. Our contributions
(A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates. This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries). See Figure 1 for an illustration. We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries [15], but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
@@
(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs. As discussed for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms. Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms. We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms:
(1) The first shows that, when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms. (2) We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets. (3) We construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private. Then imposing an upper bound on the sum of the privacy parameter $\delta_{i}$ of the created mechanisms, we introduce a new composition theorem for concurrent parallel composition with this upper bound. (4) As a special case, our new composition theorem extends the “parallel composition theorem” for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms when the composed mechanisms satisfy pure differential privacy.
(C) We formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss. These conditions are also reformulated as a set of properties for pseudocode describing continual mechanisms.
(D) We use our concurrent composition theorems for continual mechanisms, with the aforementioned general conditions, to give a simpler and more modular privacy analysis of a recent continual histogram mechanism [13]. In the case of approximate DP, ours is the first proof that the privacy of the mechanism holds against adaptive adversaries.
(E) We how that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
@@
For every fixed sequence $(\epsilon_{1},\delta_{1}),\ldots,(\epsilon_{k},\delta_{k})$ of privacy-loss parameters and verification functions $f_{1},\ldots,f_{k}$, the concurrent composition of continual mechanisms $\mathcal{M}_{i}$ that are $(\epsilon_{i},\delta_{i})$-DP w.r.t. $f_{i}$ is $(\epsilon,\delta)$-DP for the same $(\epsilon,\delta)$ that holds for composing noninteractive mechanisms that are $(\epsilon_{i},\delta_{i})$-DP.
@@
A key step in the proof of Theorem 1.2 is to show that for any $(\epsilon_{i},\delta_{i})$-DP continual mechanism $\mathcal{M}$, the interactive mechanism $\mathcal{M}^{\prime}(\cdot)=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)\!\circ^{*}\!\mathcal{M}$ can be simulated by post-processing the randomized response mechanism $\mathrm{RR}_{\epsilon_{i},\delta_{i}}(\cdot)$. Lyu [19] proves such a result under the assumptions that the answer space of $\mathcal{M}$ is finite and that there exists a predetermined upper bound $T$ on the number of interactions. We remove the latter assumption and relax the former to having an infinite (but discrete) answer space.
@@
We study the extent to which parallel composition extends to the concurrent composition of continual mechanisms, meaning the neighboring relation is as above, but the adversary can interact concurrently with the mechanisms, and in particular decide adaptively on which $k$ of them to have different update sequences. Moreover, we also consider the case where $\ell$ is unbounded, meaning the adversary (or honest analyst) can adaptively decide how many mechanisms to interact with. We refer to this as concurrent $k$-sparse parallel composition. A more restricted version of this composition, which is limited to noninteractive mechanisms, was previously studied in [20].
@@
For every $\epsilon_{0}>0$ and $0\leq\delta_{0}\leq 1$ and $k\in\mathbb{N}$, and every $(\epsilon_{0},\delta_{0})$-DP interactive mechanism $\mathcal{M}_{0}$, the concurrent $k$-sparse parallel composition of an unbounded number of copies of $\mathcal{M}_{0}$ satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms.
As discussed later in the paper, we also prove a generalization of Theorem 1.3 for the case of composing interactive mechanisms $\mathcal{M}_{i}$ with differing $(\epsilon_{i},\delta_{i})$ parameters.
@@
Theorem 1.4 (counterexample for the concurrent parallel composition of approximate DP continual mechanisms, informally stated).
@@
This theorem says that, at least in terms of the $\delta$ parameter, concurrent parallel composition of continual mechanisms does not provide any benefit over non-parallel composition, since the bound of $1-(1-\delta_{0})^{\ell}$, is the bound that we would get by just applying composition over all $\ell$ mechanisms. Inspired by this counterexample, we prove a weaker version of Theorem 1.3 for the concurrent parallel composition of continual mechanisms.
@@
Theorem 1.5 (concurrent parallel composition of approx DP continual mechanisms, informally stated).
@@
For every $\epsilon_{0}>0$, $0\leq\delta^{\prime}\leq 1$, and $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},\delta_{i})$-DP continual mechanisms, where the parameters $\delta_{i}$ are chosen adaptively such that $\sum_{i}\delta_{i}\leq\delta^{\prime}$, satisfies $(\epsilon,\delta+\delta^{\prime})$-DP if the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms is $(\epsilon,\delta)$-DP.
We note that the formal version of this result applies to $(\epsilon_{i},\delta_{i})$-DP continual mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters. The proof of Theorem 1.5 relies on a key property of our new post-processing construction (recall that $\mathrm{RR}_{\epsilon,\delta}(b)$ denotes the randomized response mechanism with parameters $\epsilon$ and $\delta$ and input $b\in\{0,1\}$):
Lemma 1.6 (post-processing of randomized response mechanisms, informal statement).
For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring datasets $x_{0},x_{1}$, there exists an interactive post-processing mechanism $\mathcal{P}$ that receives the outputs of $\mathrm{RR}_{\epsilon,0}(b)$ and $\mathrm{RR}_{0,\delta}(b)$ and simulates $\mathcal{M}(x_{b})$ identically. As long as the answer distributions of $\mathcal{M}(x_{0})$ and $\mathcal{M}(x_{1})$ to the queries asked so far are identical, $\mathcal{P}$ does not access the outcome of $\mathrm{RR}_{\epsilon,0}(b)$.
Setting $\delta^{\prime}=0$, Theorem 1.5 implies:
Corollary 1.7 (concurrent parallel composition of pure DP continual mechanisms, informally stated).
For every $\epsilon_{0}>0$ and integer $k\in\mathbb{N}$, the concurrent $k$-sparse parallel composition of $(\epsilon_{0},0)$-DP mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantees as the composition of $k$ noninteractive $(\epsilon_{0},0)$-DP mechanisms.
Like Theorem 1.5, we also prove a generalization of Corollary 1.7 for the case of composing mechanisms $\mathcal{M}_{i}$ with differing $\epsilon_{i}$ parameters.
Concurrent filter composition theorems for continual mechanisms.
Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter. Consider adversaries that create continual mechanisms with adaptive privacy parameters such that $\mathrm{Filt}(\sigma)=\top$ at every step, where $\sigma$ is the sequence of privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms $\mathcal{M}_{i}$ up to that point. Suppose each $\mathcal{M}_{i}$ is $(\epsilon_{i},\delta_{i})$-DP with respect to a verification function $f_{i}$. In addition to creating new mechanisms, these adversaries may adaptively generate pairs of inputs for existing mechanisms such that $f_{i}$ applied to the sequence of input pairs for each $\mathcal{M}_{i}$ evaluates to $\top$. The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms is said to be $(\epsilon,\delta)$-DP if the view of such adversaries is $(\epsilon,\delta)$-indistinguishable when all mechanisms receive either the first inputs from their respective pairs or all receive the second ones.
Theorem 1.8 (Concurrent filter composition for continual mechanisms).
Let $\mathrm{Filt}:\left([0,\infty)\times[0,1]\right)^{*}\to\{\top,\bot\}$ be a filter. The concurrent $\mathrm{Filt}$-filter composition of continual mechanisms satisfies the same $(\epsilon,\delta)$-DP guarantee as the $\mathrm{Filt}$-filter composition of non-interactive mechanisms.
General Conditions for Privacy Analysis of Continual Mechanisms.